The airgap in a magnetic circuit is 1.1 mm long and 2000 mm2 in cross-section. Calculate: (i) the reluctance of the airgap; and (ii) the m.m.f. to send a f lux of 700 microwebers across the airgap.
1
Expert's answer
2020-05-20T09:08:49-0400

i)


"B=\frac{\Phi}{A}=\frac{700\cdot 10^{-6}}{2000\cdot 10^{-6}}=0.35\frac{Wb}{m^2}"

ii)


"H=\frac{B}{\mu_0}=\frac{0.35}{4\pi\cdot 10^{-7}}=2.785\cdot10^5\frac{AT}{m}"